OpenStudy (rebeccapink6398):

WILL GIVE MEDAL!!!!!! PLZ O' PLZ HELP!!!!! A circle has a radius of 9 mm. What is the approximate area of the circle? A. 508.68 mm2 B. 56.52 mm2 C. 254.34 mm2 D. 1017.36 mm2

OpenStudy (rebeccapink6398):

@whpalmer4

OpenStudy (rebeccapink6398):

@Loser66

jimthompson5910 (jim_thompson5910):

Use the formula A = pi*r^2 where A= area of the circle pi = 3.14 (approximately) r = radius

jimthompson5910 (jim_thompson5910):

in this case, r = 9

OpenStudy (rebeccapink6398):

ok

jimthompson5910 (jim_thompson5910):

so A = pi*r^2 A = 3.14*9^2 A = ??

OpenStudy (rebeccapink6398):

56.52

jimthompson5910 (jim_thompson5910):

the ^2 doesn't mean "times 2"

jimthompson5910 (jim_thompson5910):

it means "squared"

OpenStudy (rebeccapink6398):

oh

OpenStudy (rebeccapink6398):

OH! sorry lol I forgot

OpenStudy (rebeccapink6398):

254.34??

jimthompson5910 (jim_thompson5910):

254.34 is correct

OpenStudy (rebeccapink6398):

yay!!!

jimthompson5910 (jim_thompson5910):

nice work